OpenStudy (anonymous):

what exponential function passes thru (0,4) and (-2,64)

OpenStudy (mattfeury):

the basic form for an exponential equation is \[y = ax^2 + b\] so I'd plug in (0,4) and we can solve for b then plug in (-2,64) to solve for a

OpenStudy (mattfeury):

\[4 = 0*0^2 + b\]\[b = 4\]\[64 = a*(-2)^2 + 4\]\[15 = a\] boom. plug those in your general equation (above) and you've got an answer!

OpenStudy (anonymous):

these are the choices tho 1) f(x)=4(1\4)^x 2) f(x)=4^x 3)f(x)=4(1\4)^-x 4)f(x)=-4(1\4)^x

OpenStudy (anonymous):

thats why im confused?

OpenStudy (mattfeury):

well you can plug each set of numbers into all those functions and see which works. it looks like (1) will work for both sets.

OpenStudy (mattfeury):

ahh. i misread 'exponential' to be 'quadratic'. My answer is one example, but there are many functions that would work.

OpenStudy (anonymous):

thank you so much and by any chance do you know this one: 3 sides of a rectangular garden are enclosed using 120 meters of fencing. what is the maximum area of garden?

OpenStudy (mattfeury):

hmm what level math are you in? I am tempted to use derivatives, but not if you haven't learned them yet.

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

and yeah i havent learned that yet lol

OpenStudy (mattfeury):

well I'm thinking that the maximum area would be a perfect square rather than a rectangle, don't you think? this would mean each of the three sides are equal.

OpenStudy (anonymous):

The basic form is\[y=a^{bx}\]

myininaya (myininaya):

ok we have x+x+y=120 so y=120-2x so A=xy=x(120-2x)=120x-2x^2 so if we find the vertex of this parabola we can find the max

OpenStudy (anonymous):

I get \[y=2^{-3x}\]

myininaya (myininaya):

i got y=4^(-3x/2) for that one

OpenStudy (anonymous):

Actually, I don't...!

OpenStudy (anonymous):

so with the choices 1) 3,600m^2 2) 1,800m^2 3)1,600m^2 4)900m^2 whats the correct answer

myininaya (myininaya):

anyways the vertex of that parabola is (0,120) and you find that by writing the parabole as a*(x-h)^2+k where (h,k) is the vertex

myininaya (myininaya):

oops one sec i made a mistake

OpenStudy (anonymous):

\[y=4e^{-x \ln4}\]

OpenStudy (anonymous):

If you let y be length of one of the sides, and x be the length of the other two sides, then y+2x=120. The area is given by A=xy, so since y=120-2x, just substitute for y in A=xy. This gives A=x(120-2x). You are trying to find x that maximizes this. Since the graph of this will be an upside down parabola with roots at x=0 and x=60, this means the maximum will occur at x=30 (halfway between the roots). So the max area is A=30*(120-2*30)=30*60=1800.

OpenStudy (anonymous):

When x is 0, y is 4. When x is -2, y is 64.

myininaya (myininaya):

the vertex of the parabola is (30,1800) since A=-2(x-30)^2+1800

OpenStudy (anonymous):

Why are we talking about quadratics?

OpenStudy (anonymous):

oh okay so the answe is 2

OpenStudy (anonymous):

1800

OpenStudy (anonymous):

\[y=ae^{bx}\]When x=0, y=4, so\[4=ae^0=a\]so a=4. When x=-2, y=64, so\[64=4e^{-2b} \rightarrow 16=e^{-2b} \rightarrow \ln 16 = -2b \rightarrow b=-\frac{1}{2}\ln 16=-\ln 4\]so b = -ln(4). hence your equation is,\[y=4e^{-(\ln 4)x }\]

myininaya (myininaya):

you're so smart lokisan lol

OpenStudy (anonymous):

haha...but wait, there's more...

OpenStudy (anonymous):

\[e^{-(\ln4)x}=e^{\ln \frac{1}{4}x}=\left( e^{\ln \frac{1}{4}} \right)^x=\frac{1}{4^x}\]so\[y=4e^{-(\ln 4)x}=4.\frac{1}{4^x}=\frac{1}{4^{x-1}}=4^{1-x}\]

OpenStudy (anonymous):

So your equation can be written as,\[y=4^{1-x}\]

OpenStudy (anonymous):

Silence...

OpenStudy (anonymous):

lol u guys are so good at this im so confused. but okay if thats my equation whats the answer, 1 2 3 or 4

OpenStudy (anonymous):

3600 or 1800 or 1600 or 900

OpenStudy (anonymous):

What do you actually need to find out? I thought you just had to come up with the equation.

OpenStudy (anonymous):

imrickjames is here...haven't seen you in a while ;]

OpenStudy (anonymous):

noo .. i wrote up there earlier the choices...its asking me what the maximum area of the garden is

OpenStudy (anonymous):

Okay...let me check the multiple choices...

OpenStudy (anonymous):

What garden? Isn't that another question?

OpenStudy (anonymous):

Oh, I see the garden question. So the others didn't work it out?

OpenStudy (anonymous):

no i just need to answer the garden question. i need to choose one of those 4 and show work.

OpenStudy (anonymous):

1800

OpenStudy (anonymous):

Thanks for fanning me, Matt :)

OpenStudy (anonymous):

thank u ..wat about if a car is purchased for 20,000. the value of the car decreases by 10% each year. about how many years will it take for the vallue of the car to reach 8,000

OpenStudy (mattfeury):

hey! you can't know it was me! heh :). impressed by your scan.

OpenStudy (anonymous):

About 8.7 years.

OpenStudy (anonymous):

Impressed by the scan...so *that's* how I increase the fan base ;)

OpenStudy (anonymous):

how did u get that thoo

OpenStudy (anonymous):

These types of equations with growth and decay in proportion to themselves follow the form\[p(t)=p(0)e^{kt}\]where p(0) is the initial value and k some constant. For growth, it's positive, for decay, negative. You get this equation by solving the differential equation,\[\frac{dp}{dt}=kp\]which is separable (I don't know what level you're at). Anyway, with this equation, you use your data to find out what p(0) and k are, then answer the question.

OpenStudy (anonymous):

Here, p(0) is 20,000. To find k, you know that after 1 year, the car's value is 10% less, so it's value at time t=1 is 18,000. So,\[18,000=20,000e^{k}\]Divide both sides by 20,000 and take the natural log to find k. Now you have your p(0) and your k, so you can find p(t) at any time t.

OpenStudy (anonymous):

okay thank uuu ! and if i have to find |3+i| that equals what ?

OpenStudy (anonymous):

You want to find t when p(t) is 8000, so you set it up as,\[8000=20000e^{ \ln (\frac{9}{10})t}\]You have to solve for t. Divide both sides by 20,000 and again, take the natural log to find\[\frac{8000}{20000}=\ln (\frac{9}{10})t \rightarrow t =\frac{8000/20000}{\ln (9/10)} \approx 8.7yrs\]

OpenStudy (anonymous):

It's asking for the magnitude. The magnitude of a complex number is z is,\[|z|=|x+iy|=\sqrt{x^2+y^2}\]You consider the real and imaginary part (note, the imaginary part is NOT iy, it's y, and y is real...yi is imaginary).

OpenStudy (anonymous):

\[|3+i|=\sqrt{3^2+1^2}=\sqrt{10}\]

OpenStudy (anonymous):

Good luck. If you have anymore questions, post them separately...I need to log out :D

OpenStudy (anonymous):

thank youuuuu so muchhh